OpenStudy (turingtest):

Last integral of the night for me\[\int e^{4x}\sqrt{1+e^{2x}}dx\]I just think it's fun.

myininaya (myininaya):

hint: \[\int\limits_{}^{}2 \cdot \frac{1}{2} e^{2x} e^{2x} \sqrt{1+e^{2x}} dx\]

OpenStudy (mr.math):

\(e^{2x}=\tan^2(u)\) would work, I think.

OpenStudy (kinggeorge):

without solving, my first instinct would be to do a u-sub for \[e^{2x}\]

myininaya (myininaya):

omg why mr.math

myininaya (myininaya):

don't make it hard

myininaya (myininaya):

u=1+e^(2x)

myininaya (myininaya):

\[u=1+e^{2x} => du= 2 e^{2x} dx ; u=1+e^{2x} => u-1=e^{2x}\]

myininaya (myininaya):

\[\int\limits\limits_{}^{}2 \cdot \frac{1}{2} e^{2x} e^{2x} \sqrt{1+e^{2x}} dx \] \[\frac{1}{2}\int\limits_{}^{} (u-1) \sqrt{u} du\]

OpenStudy (turingtest):

I did it Mr.Math's way. Glad I asked, this seems much simpler. G'nigh y'all :D

myininaya (myininaya):

lol

myininaya (myininaya):

boys are cute

myininaya (myininaya):

;)

OpenStudy (turingtest):

:)